# Persistence of AR($1$) sequences with Rademacher innovations and linear mod $1$ transforms

### v2

@@ 2.3. The aggregated chain approach
In this section we explain the approach based on reduction to a Markov chain obtained by lumping the states between the points of the orbit of 00 under $T_{a}$. This Markov chain has a finite number of states when $\varkappa^{\prime}_{a}<\infty$, which allows one to find the asymptotics in the exit problem using the classical Perron–Frobenius theorem for non-negative matrices.
@@
Corollary 6.
For any $x\in[0,\frac{1}{1-a}]$, the sequence $f(X_{n})\mathbbm{1}\{\tau>n\}$ is a time-homogeneous non-negative Markov chain under $\mathbb{P}_{x}$ and it is true that $\tau=\inf\{n\geq 1:f(X_{n})=0\}$. If $\varkappa^{\prime}_{a}<\infty$, this chain has $\varkappa^{\prime}_{a}+2$ states, including the absorbing state at zero.
The latter equalities follow from
$T_{a}^{k-1}(0)=\begin{cases}aT_{a}^{k}(0)-1,\quad\text{if }T_{a}^{k-1}(0)<1,\\ aT_{a}^{k}(0)+1,\quad\text{if }T_{a}^{k-1}(0)\geq 1.\end{cases}$
Examples: Suppose that $\varkappa^{\prime}_{a}=2$. Equivalently,
$\frac{2a-1}{1-a}<T_{a}^{2}(0)=\frac{1-a}{a^{2}}\leq 1,$
that is
[…]
In this case the transition matrix of the lumped chain is given by
i.e. approximately, $0.58250.5437\leq a<$ 0.5825. In theses cases the transition matrices of the lumped chains are given respectively by
$\begin{pmatrix}1&0&0&0&0\\ 1-p&0&0&p&0\\ 0&1-p&p&0&0\\ 1-p&0&p&0&0\\ 0&0&0&1-p&p\\ \end{pmatrix}\text{ and }\begin{pmatrix}1&0&0&0&0\\ 1-p&0&0&0&p\\ 0&1-p&p&0&0\\ 1-p&0&p&0&0\\ 1-p&0&0&p&0\\ \end{pmatrix}.$
Remark 7.
We do not know how to make use of the lumped Markov chain $f(X_{n})\mathbbm{1}\{\tau>n\}$ when $\varkappa^{\prime}_{a}=\infty$. In particular, it is unclear how to describe concisely the transition probabilities of this chain because it may have uncountably many states even though every function $f\in U_{+}$ has countably many jumps. $\diamond$
Proof of Corollary 6.
We first restrict both function $f$ and the stopped Markov chain $\{X_{n\wedge\tau}\}$ to the set $(-\infty,\frac{1}{1-a}]$. Then Proposition 5 implies that the distribution of $f(X_{1\wedge\tau})$ under $\mathbb{P}_{z}$ does not change as $z$ varies over any level set of $f$. This condition is known to ensure, see e.g. [10, Corollary 1] (this particular reference applies only for countable chains), that the lumped sequence $\{f(X_{n\wedge\tau})\}_{n\geq 0}$ is a Markov chain on the range of $f$. This set has cardinality $\varkappa_{a}^{\prime}+2$ when $\varkappa_{a}^{\prime}<\infty$. The state zero is clearly absorbing. Since $f=0$ on $(-\infty,0)$, we have $f(X_{n\wedge\tau})=f(X_{n})\mathbbm{1}\{\tau>n\}$. Thus, $f(X_{n})\mathbbm{1}\{\tau>n\}$ is a Markov chain. The equality for $\tau$ follows because $f>0$ on $[0,\infty)$ by $f\in U_{+}$. ∎
Proof of Proposition 5.
Assume that $a\in(0,\frac{1}{2})$. We have $\varkappa_{a}^{\prime}=0$ and thus $f$ is constant on $[0,\frac{1}{1-a}]$. Therefore, $f(X_{n\wedge\tau_{x}}^{x})$ stays constant until $\{X_{n}\}$ leaves $[0,\frac{1}{1-a}]$ at the moment $\tau_{x}$, which is the same for all $x$ by $\tau_{x}=\inf\{n\geq 1:\xi_{n}=-1\}$. This yields the claims.
Assume that $a\in[\frac{1}{2},\frac{2}{3})$. From the representation
@@
Denote by $r$ the right endpoint of $J$. By right-continuity of $f$, the value of $f$ at the left endpoint of $aJ-1$ equals the value of $f$ on $\operatorname*{int}(aJ-1)$, and the same applies to $aJ+1$. This establishes constancy of $f$ on the whole of $aJ-1$ and on the whole of $aJ+1$ if $r\not\in J$ because in this case the intervals $aJ\pm 1$ do not include their right endpoints.
@@
Lastly, the equality $\tau_{x}=\tau_{y}$ follows from (21) and the facts that $f=0$ on $(-\infty,0)$ and $f>0$ on $[0,\infty)$. ∎
@@ 3. Existence of a positive solution to equation (11)
which is the l.h.s. of (11). We need to show that the equation $R_{a}(\lambda)=1$ has exactly one positive solution $\lambda=\lambda_{a}$. This equation will arise naturally in the next section. We will establish existence together with the following bounds.
@@
Let $a\in(\frac{1}{2},\frac{2}{3}]$. Then equation (11) has a unique positive solution $\lambda_{a}$. Moreover, there exists a constant $C>0$ such that for every $p\in(0,1)$, we have
@@
for all $x\in[0,1/(1-a)]$ and all integers $0\leq k\leq n\leq\varkappa_{a}(x)$. The first inequality in (22) is strict unless $a=\frac{2}{3}$ and $p=\frac{1}{2}$.
@@
It takes us much more effort to prove Proposition 8 when $\varkappa_{a}=\infty$ and $q>p$. The key to our proof is the following assertion, where we essentially compare the (inverse) number of 00’s in the first digits of $(\frac{1}{a},\frac{1}{2})$-expansions of 00 and of an arbitrary point $x$. This is in the spirit of Theorem 1 by Parry [18]. To state our result, denote by $\sigma_{a}(x)$ the total number of returns to $[0,1)$ of the trajectory of $x$ under $T_{a}$ and by $t_{k}(x)$ the corresponding return times:
@@
$d_{n}=\begin{cases}t_{1},&n=0,\\ \min\big{\{}\frac{t_{1}+1}{1},\ldots,\frac{t_{n}+1}{n},\frac{t_{n+1}}{n+1}\big{\}},&1\leq n<\sigma_{a},\\ \min\big{\{}\frac{t_{1}+1}{1},\ldots,\frac{t_{n}+1}{n}\big{\}},&n=\sigma_{a}\text{ and }\sigma_{a}<\infty.\\ \end{cases}$
@@
$T_{a}^{t_{i}}(0)\leq T_{a}^{t_{i}}(x)=T_{a}^{t_{i}(x)}(x)<\frac{2a-1}{1-a},\qquad 0\leq i\leq\min(m,\sigma_{a}-1)=:m^{\prime}.$ (25)
We prove this by induction, assuming that $n\geq 1$, otherwise there is noting to prove. In the basis case $i=0$ this is true by the assumption. If we already established this for all $i\leq j$ for some $j<m^{\prime}$, then $1\leq T_{a}^{t_{j}+1}(0)\leq T_{a}^{t_{j}+1}(x)$ since $T_{a}$ increases on the interval $[0,\frac{2a-1}{1-a})$. Then $T_{a}^{\ell}(0)\leq T_{a}^{\ell}(x)$ for all $t_{j}+1\leq\ell\leq t_{j+1}$ since $T_{a}$ increases on $[1,\frac{1}{1-a}]$. Moreover, $T_{a}^{t_{j+1}}(x)=T_{a}^{t_{j+1}(x)}(x)<\frac{2a-1}{1-a}$ since $j+1\leq m^{\prime}<\min(\sigma_{a}(x),\sigma_{a})$. This establishes (25).
@@
[…]
@@
a) We argue by induction. In the basis case $k=1$ we have $D_{1}=\{\frac{2a-1}{1-a},\frac{1}{1-a}\}$, hence the set $\{x:\varkappa_{a}(x)\geq 1\}$, which is $[0,\frac{1}{1-a}]\setminus I_{a}$, is of the form stated. The claims concerning $\delta_{0}(x)$ and $T_{a}(x)$ are clearly satisfied.
@@
If $\frac{2a-1}{1-a}<T_{a}^{n}(g_{n}(y))<1$, then by $T_{a}^{n}(y)=\frac{1}{1-a}>1$, there is a unique $z\in[g_{n}(y),y]$ such that $T_{a}^{n}(z)=1$. Then $z\in G_{n+1}$ by $T_{a}^{n+1}(z)=T_{a}(1)=0$, and $y\in D_{n+1}$. By the same argument as in the previous case, $[z,y]$ is the maximal subinterval of $[g_{n}(y),y]$ contained in $\text{dom}(T^{n+1}_{a})$, hence $g_{n+1}(y)=z$.
Lastly, if $0\leq T_{a}^{n}(g_{n}(y))\leq\frac{2a-1}{1-a}$, then by $T_{a}^{n}(y)>1$, there exist unique $z_{1},z_{2}\in[g_{n}(y),y]$ such that $T_{a}^{n}(z_{1})=\frac{2a-1}{1-a}$ and $T_{a}^{n}(z_{2})=1$. Similarly to the previous cases, $[g_{n}(y),z_{1}]$ and $[z_{2},y]$ are the maximal subintervals of $[g_{n}(y),y]$ contained in $\text{dom}(T^{n+1}_{a})$, and there are no other ones. It holds $z_{1},y\in D_{n+1}$ and $g_{n}(y),z_{2}\in D_{n+1}$. We also have $g_{n+1}(z_{1})=g_{n}(y)$, $g_{n+1}(y)=z_{2}$.
@@
Furthermore, in either of the three cases, the function $T_{a}^{n+1}$ is continuous and strictly increasing on $[g_{n+1}(y),y]$ as a composition of $T_{a}$ and $T_{a}^{n}$, which have these properties on $[1,\frac{1}{1-a}]$ and $[g_{n}(y),y]$, respectively. Then $\delta_{n}$ is constant on $[g_{n+1}(y),y]$. In the third case, $T_{a}^{n+1}$ and $\delta_{n}$ also have these respective properties on $[g_{n+1}(z_{1}),z_{1}]$ because on this interval $T_{a}^{n}$ does not exceed $\frac{2a-1}{1-a}$ and $T_{a}$ is continuous and strictly increasing on $[0,\frac{2a-1}{1-a}]$.
b) The proof for $a=\frac{2}{3}$ is analogous. It suffices to replace throughout $[g_{n}(y),y]$ by $[g_{n}(y-),y)$ for all $y\in D_{n}\setminus\{3\}$ and use that $T_{2/3}^{n}(y-)=3$ instead of $T^{n}_{a}(y)=\frac{1}{1-a}$. ∎
@@
because the l.h.s. is integer. This proves inequality (23) when $\sigma_{a}(x)=\infty$.
@@
We first assume that $g_{n}(y)\neq y$. Since the functions $z\mapsto L_{k}(z)$ for $0\leq k\leq n$ are constant on $[g_{n}(y),y]$ by Lemma 10.a, inequality (23) follows instantly from the result in the previous case if we prove that there is an $x^{\prime}\in[g_{n}(y),y]$ such that $\sigma_{a}(x^{\prime})=\infty$. We will use that there exist periodic orbits of $T_{a}$ starting arbitrarily close to the left of $y$. Consider two cases. If $T_{a}^{m}(y)=\frac{2a-1}{1-a}$ for a unique $10\leq m\leq n-$ 1, then $T_{a}^{m}<\frac{2a-1}{1-a}$ on $[g_{n}(y),y)$ by Lemma 10.a. By piecewise continuity of $T_{a}$, the equation $z=T_{a}^{t_{1}(z)}(z)$ admits infinitely many solutions, which accumulate to $\frac{2a-1}{1-a}$ from the left. Pick any solution $z^{\prime}\in[T_{a}^{m}(g_{n}(y)),\frac{2a-1}{1-a})$, and then take the unique $x^{\prime}\in[g_{n}(y),y]$ that solves $T_{a}^{m}(x^{\prime})=z^{\prime}$. In the second case when $y=\frac{1}{1-a}$, we argue similarly and take any $x^{\prime}\geq g_{n}(y)$ that satisfies $x^{\prime}=T_{a}^{t_{1}(x^{\prime})+1}(x^{\prime})$.
We now assume that $g_{n}(y)=y$. This is possible only when $G_{n}$ intersects $D_{n}$, that is when the orbit of zero hits $\frac{2a-1}{1-a}$. Then $T_{a}^{k_{1}}(x)=0$ and $T_{a}^{k_{2}}(x)=\frac{2a-1}{1-a}$ for the unique $k_{1}$, $k_{2}$ such that $10\leq k_{1}<k_{2}\leq n-$ 1. Thus, $x\in G_{k_{2}}$ and by Lemma 10.a, there exists a unique $y^{\prime}\in D_{k_{2}}$ such that $g_{k_{2}}(y^{\prime})=x$. In other words, $\varkappa_{a}(z)=k_{2}$ for every $z\in(g_{k_{2}}(y^{\prime}),y^{\prime})$. Clearly, $g_{k_{2}}(y^{\prime})\neq y^{\prime}$, and we can apply the result of the previous case for all $k$ satisfying $0\leq k\leq k_{2}$. Therefore, since the functions $z\mapsto L_{k}(z)$ for $0\leq k\leq k_{2}$ are constant on $[g_{k_{2}}(y^{\prime}),y^{\prime}]$,

### v3

@@ 2.2. Reduction to a finite chain when $\varkappa^{\prime}_{a}<\infty$
In this section we explain the approach based on reduction to a Markov chain obtained by lumping the states between the points of the orbit of 0 under $T_{a}$. This Markov chain has a finite number of states when $\varkappa^{\prime}_{a}<\infty$, which allows one to find the asymptotics in the exit problem using the classical Perron–Frobenius theorem for non-negative matrices. This also allows us to determine the exponent $\lambda_{a}$ numerically.
@@
Proof.
Assume that $a\in(0,\frac{1}{2})$. We have $\varkappa_{a}^{\prime}=0$ and thus $f$ is constant on $0,\frac{1}{1-a}]$. Therefore, $f(X_{n\wedge\tau_{x}}^{x})$ stays constant until $\{X_{n}\}$ leaves $0,\frac{1}{1-a}]$ at the moment $\tau_{x}$. Noting now that $\tau_{x}=\inf\{n\geq 1:\xi_{n}=-1\}$ for every $x\in[0,\frac{1}{1-a}]$, we obtain the equalities $\tau_{x}=\tau_{y}$ and (25).
Assume now that $a\in[\frac{1}{2},\frac{2}{3})$. From the representation
@@
Denote by $r$ the right endpoint of $J$. By right-continuity of $f$, the value of $f$ at the left endpoint of aJ1aJ-1 equals the value of $f$ on $\operatorname*{int}(aJ-1)$, and the same applies to $aJ+1$. This establishes constancy of $f$ on the whole of aJ1aJ-1 and on the whole of $aJ+1$ if $r\not\in J$ because in this case the intervals $aJ\pm 1$ do not include their right endpoints.
@@
Lastly, the equality $\tau_{x}=\tau_{y}$ follows from (25) and the facts that $f=0$ on $(-\infty,0)$ and $f>0$ on $0,\infty)$. ∎
Corollary 7.
For any $x\in[0,\frac{1}{1-a}]$, the sequence $f(X^{x}_{n})\mathbbm{1}\{\tau_{x}>n\}$ is a time-homogeneous non-negative Markov chain under $\mathbb{P}_{x}$ and it is true that $\tau_{x}=\inf\{n\geq 1:f(X_{n})=0\}$. If $\varkappa^{\prime}_{a}<\infty$, this chain has $\varkappa^{\prime}_{a}+2$ states, including the absorbing state at zero.
Proof.
We first restrict both function $f$ and the stopped Markov chain $\{X_{n\wedge\tau}\}$ to the set $(-\infty,\frac{1}{1-a}]$. Then Proposition 6 implies that the distribution of $f(X_{1\wedge\tau})$ under $\mathbb{P}_{z}$ does not change as $z$ varies over any level set of $f$. This condition is known to ensure, see e.g. [11, Corollary 1] (this particular reference applies only for countable chains), that the lumped sequence $\{f(X_{n\wedge\tau})\}_{n\geq 0}$ is a Markov chain on the range of $f$. This set has cardinality $\varkappa_{a}^{\prime}+2$ when $\varkappa_{a}^{\prime}<\infty$. The state zero is clearly absorbing. Since $f=0$ on $(-\infty,0)$, we have $f(X_{n\wedge\tau})=f(X_{n})\mathbbm{1}\{\tau>n\}$. Thus, $f(X_{n})\mathbbm{1}\{\tau>n\}$ is a Markov chain. The equality for $\tau$ follows because $f>0$ on $0,\infty)$ by $f\in U_{+}$. ∎
The latter equalities follow from
[…]
Let us give two examples. Suppose that $\varkappa^{\prime}_{a}=2$. Equivalently,
$\frac{2a-1}{1-a}<T_{a}^{2}(0)=\frac{1-a}{a^{2}}\leq 1,$
that is
$0.618\approx\frac{\sqrt{5}-1}{2}\leq a<\frac{1}{6}\left(2-\sqrt[3]{6\sqrt{33}-26}+\sqrt[3]{6\sqrt{33}+26}\right)\approx 0.6478.$
In this case the transition matrix of the lumped chain is given by
i.e. approximately, $0.5437\leq a<0.5825$. In theses cases the transition matrices of the lumped chains are given respectively by
[…]
Remark 8.
We do not know how to make use of the lumped Markov chain $f(X_{n})\mathbbm{1}\{\tau>n\}$ when $\varkappa^{\prime}_{a}=\infty$. In particular, it is unclear how to describe concisely the transition probabilities of this chain because it may have uncountably many states even though every function $f\in U_{+}$ has countably many jumps.
In the next section we present a different approach, which works for all values of $a\in(0,\frac{2}{3}]$, unlike the reduction described above. This method is based on the compactness properties of the transition operator of the killed chain $\{X_{n}\}$.
2.3. Compactness approach
Let $P$ be the transition operator of $\{X_{n}\}$ killed at exiting $0,\frac{1}{1-a}]$. It acts on a measurable bounded function $f$ on $0,\frac{1}{1-a}]$ as follows:
$Pf(x)=\mathbb{E}_{x}[f(X_{1});\tau>1].$
Under assumption (4) that the innovations are $\pm 1$, we have
$Pf(x)=pf(ax+1)+qf(ax-1)\mathbbm{1}\{x\geq 1/a\}.$ (26)
Since
$\mathbb{P}_{x}(\tau>n)=P^{n}\mathbbm{1}(x),$
we seek to find a Banach space of functions on [0,1/(1a)][0,1/(1-a)] where $P$ would be compact or quasi-compact, expecting that the asymptotics of $P^{n}\mathbbm{1}$ is defined by the spectral radius of $P$, which should be the largest eigenvalue of $P$. To this end, let us consider the behaviour of the set of discontinuities $D_{f}$ of a function $f$ under the action of $P$, with continuity at the endpoints 0 and 1/(1a)1/(1-a) understood as one-sided continuity.
1. The case $a\in(0,\frac{2}{3})$.
Assume that $a\in[\frac{1}{2},\frac{2}{3})$. If $Pf$ is discontinuous at an $x\in[0,\frac{1}{1-a}]$, then the same holds true for at least one of the two terms in (26). Note that $f(ax+1)$ is discontinuous at $x$ if and only if $f$ is discontinuous at $y=ax+1$; here $y\in[1,\frac{1}{1-a}]$ and $x=\frac{1}{a}(y-1)$. On the other hand, $f(ax-1)\mathbbm{1}\{x\geq 1/a\}$ is discontinuous at $x$ if and only if $f$ is discontinuous at y=ax1y=ax-1 or it holds that $x=1/a,f(0)\neq 0$; here $y\in[0,\frac{2a-1}{1-a}]$ and $x=\frac{1}{a}(y+1)$ in both cases.
The observations above can be summed up using the mapping
$D_{Pf}\subset T_{a}\big((D_{f}\cup\{0\})\setminus I_{a}\big).$ (27)
It follows from (27) that for $a\in(0,\frac{2}{3})$, the set of measurable bounded functions on $0,\frac{1}{1-a}]$ that are continuous at every point outside of $\{T_{a}^{k}(0):0\leq k<\varkappa_{a}+1\}$ is closed under the action of $P$; recall that
$\varkappa_{a}=\inf\{k\geq 0:\,T_{a}^{k}(0)\in I_{a}\}.$
This suggests to consider the set of functions
on $0,\frac{1}{1-a}]$, because we know how $P$ transforms the jumps of functions and every function in $U$ is defined by its jumps. The idea to consider general right-continuous step functions with countably many jumps (called saltus functions) in the context of linear mod 1 transforms goes back to Halfin [16].
We will show that the set $U$ is closed under the action of $P$. Then we will give a simple, explicit description of this action and show that $P$ is a quasi-compact operator on $U$; see Proposition 14 and Remark 15.
2. The case $a\in[\frac{2}{3},1)$.
Here $\frac{2a-1}{1-a}\geq 1$ and thus the mapping $T_{a}$ does not account for the discontinuities of $Pf$ arising from the discontinuities of $f$ on $1,\frac{2a-1}{1-a}]$ due to the second term in (26). The argument above gives
$D_{Pf}\subset T_{a}\big(D_{f}\cup\{0\}\big)\cup\left\{\frac{1}{a}(y+1):y\in D_{f}\cap\Big[1,\frac{2a-1}{1-a}\Big]\right\}.$ (29)
The right-hand side simplifies when $f$ has no discontinuities in $1,\frac{2a-1}{1-a}]$. This motivates us to introduce
$\varkappa_{a}=\inf\left\{k\geq 0:\,T_{a}^{k}(0)\in\Big[1,\frac{2a-1}{1-a}\Big]\right\}.$
which holds true for every $a\in[\frac{1}{2},\frac{2}{3}]$ and follows by simple induction. Hence if $T_{2/3}^{k}(0)=1$, then $2^{k}=-\sum_{i=0}^{k-1}(-1)^{\delta_{i}}3^{k-i}2^{i}$, which is impossible because the right-hand side is odd.
With this extended definition of $\varkappa_{a}$, the functional space $U$ introduced in (28) is again closed under the action of $P$ if $a\in[\frac{2}{3},1)$ is such that $\varkappa_{a}=\infty$. The set of such $a$ is contained in $\frac{2}{3},\frac{\sqrt{2}}{2})$ because $T_{a}(0)\not\in[1,\frac{2a-1}{1-a}]$ only when $\frac{2a-1}{1-a}<\frac{1}{a}$, hence $2a^{2}<1$. It is possible to show that this set has Lebesgue measure zero. Note that it contains points other than $a=2/3$, for example, the unique solution to $T_{a}^{5}(0)=1/a$ on $[2/3,1)$, which is $a\approx 0.691$.
Remark 9.
Our method of proving tail asymptotics (14) seem to work unchanged for every $a\in(\frac{2}{3},1)$ such that $\varkappa_{a}=\infty$.
2.4. Transfer operator approach
Assume that $a=\frac{2}{3}$, and use the shorthand $T=T_{2/3}$. The mapping $T$ is defined on the whole of $0,\frac{1}{1-a}]$. Equation (30) now means that the killed transition operator $P$ is a weighted transfer (or the Ruelle) operator associated with $T$, where the weight takes two values $p$ and $q$. The weighted transfer operators are considered e.g. in the book by Baladi [5]. The important particular case is the standard transfer (or the Perron–Frobenius) operator $P_{T}$, defined by $P_{T}f(x)=\sum_{y\in T^{-1}(x)}f(y)/|T^{\prime}(y)|$ for $x\in[0,3]$. If $f\geq 0$, then $P_{T}f$ is the density of the measure with density $f$ pushed forward by $T$.
Assume now that $p=\frac{1}{2}$. Then by (30), we have
$Pf(x)=\frac{3}{4}P_{T}f(x),\qquad 0\leq x<3,$
and thus $P=\frac{3}{4}P_{T}$ as operators on $L^{1}([0,3])$. Since $T$ is piecewise linear and expanding (i.e. $\operatorname*{ess\,inf}_{0\leq x\leq 3}|T^{\prime}(x)|>1$), the operator $P_{T}$ is quasi-compact on the space of functions of bounded variation on $[0,3]$ (with a.e. equal functions identified); see e.g. Boyarsky and Góra [7, Theorem 7.2.1]. The leading eigenvalue of $P_{T}$ is simple and equals $1$, and this gives $\lambda_{2/3}(1/2)=3/4$ and also implies (14), together with an alternative way of finding the function $V$. To explain this, note that the eigenfunction $h$ corresponding to the leading eigenvalue $1$ of $P_{T}$ is the invariant density under the transformation $T$. Its scaled version $\widehat{h}(x)=3h(3x)$ on $[0,1]$ is invariant under $\widehat{T}_{2/3}$. This density was found explicitly by Parry [21], and it is given (up to a multiplicative factor) by formula (68) below. Simplifying this formula and rescaling back to $[0,3]$, we can recover our function $V$ given in (16) for $a=\frac{2}{3}$ and $p=\frac{1}{2}$. In this case $h=cV/3$ is a probability density. This normalization can be seen by integrating $V$ and combining formula (40) for $c$ with the equations $\frac{1}{3}T^{k}_{2/3}(0)=1-\sum_{n\geq 0}^{\infty}(\frac{2}{3})^{n+1}\delta_{k+n}$ for integer $k\geq 0$, which in turn follow from (10) and (11).
We also have an alternative proof of our result that the quasi-invariant distribution $\nu_{2/3}$ is uniform when $p=\frac{1}{2}$. Indeed, it is easy to show that the density of $\nu_{2/3}$ is an eigenfunction of the composition (or the Koopman) operator that is dual to $P_{T}$. The constant density is its eigenfunction and there are no other eigenfunctions by [7, Theorem 3.5.2]. Therefore, the density of $\nu_{2/3}$ is constant.
@@ 3. Existence of a positive solution to equation (15)
which represents the left-hand side of (15). We need to show that the equation $R_{a}(\lambda)=1$ has exactly one positive solution $\lambda=\lambda_{a}$. This equation arises naturally in the next section. We will establish existence together with the following bounds.
@@
Let $a\in(\frac{1}{2},\frac{2}{3}]$. Then for every $p\in(0,1)$, equation (15) has a unique positive solution $\lambda_{a}=\lambda_{a}(p)$. Moreover, there exists a constant $C=C_{a}>0$ such that for every $p\in(0,1)$, we have
@@
for all $x\in[0,1/(1-a)]$ and all integers $0\leq k\leq n\leq\varkappa_{a}(x)$. The first inequality in (31) is strict unless $a=\frac{2}{3}$ and $p=\frac{1}{2}$, in which case $\lambda_{2/3}(1/2)=3/4$.
@@
It takes us much more effort to prove Proposition 10 when $\varkappa_{a}=\infty$ and $q>p$. The key to our proof is the following assertion, where we essentially compare the (inverse) number of 0’s among the first digits of the $(\frac{1}{a},\frac{1}{2})$-expansions of 0 and of an arbitrary $x\in[0,1)$. This is in the spirit of Theorem 1 by Parry [20]. To state our result, denote by $\sigma_{a}(x)$ the total number of returns to $[0,1)$ of the trajectory of $x$ under $T_{a}$ and by $t_{k}(x)$ the corresponding return times:
@@
[…]
@@
$T_{a}^{t_{i}}(0)\leq T_{a}^{t_{i}}(x)=T_{a}^{t_{i}(x)}(x)<\frac{2a-1}{1-a},\qquad 0\leq i\leq m^{\prime}.$ (34)
We prove this by induction, assuming that $n\geq 1$, otherwise there is noting to prove. In the basis case $i=0$ this is true by the assumption. If we already established this for all $i\leq j$ for some $j<m^{\prime}$, then $1\leq T_{a}^{t_{j}+1}(0)\leq T_{a}^{t_{j}+1}(x)$ since $T_{a}$ increases on the interval $0,\frac{2a-1}{1-a})$. Then $T_{a}^{\ell}(0)\leq T_{a}^{\ell}(x)$ for all $t_{j}+1\leq\ell\leq t_{j+1}$ since $T_{a}$ increases on $1,\frac{1}{1-a}]$. Moreover, $T_{a}^{t_{j+1}}(x)=T_{a}^{t_{j+1}(x)}(x)<\frac{2a-1}{1-a}$ since $j+1\leq m^{\prime}<\min(\sigma_{a}(x),\sigma_{a})$. This establishes (34).
@@
$G_{k}=\bigcup_{n=0}^{k}T_{a}^{-n}(0)\cap\{x:\varkappa_{a}(x)\geq k\}\quad\text{and}\quad D_{k}=\bigcup_{n=0}^{k-1}T_{a}^{-n}\left(\frac{2a-1}{1-a}\right)\cup\left\{\frac{1}{1-a}\right\},$
@@
a) We argue by induction. In the basis case $k=1$ we have $D_{1}=\{\frac{2a-1}{1-a},\frac{1}{1-a}\}$, hence the set $\{x:\varkappa_{a}(x)\geq 1\}$, which is $0,\frac{1}{1-a}]\setminus I_{a}$, is of the form stated. The claims concerning $\delta_{0}(x)$ and $T_{a}(x)$ are clearly satisfied.
@@
If $\frac{2a-1}{1-a}<T_{a}^{n}(g_{n}(y))<1$, then by $T_{a}^{n}(y)=\frac{1}{1-a}>1$, there is a unique $z\in[g_{n}(y),y]$ such that $T_{a}^{n}(z)=1$. Then $z\in G_{n+1}$ by $T_{a}^{n+1}(z)=T_{a}(1)=0$, and $y\in D_{n+1}$. By the same argument as in the previous case, $[z,y]$ is the maximal subinterval of $g_{n}(y),y]$ contained in $\text{dom}(T^{n+1}_{a})$, hence $g_{n+1}(y)=z$.
Lastly, if $0\leq T_{a}^{n}(g_{n}(y))\leq\frac{2a-1}{1-a}$, then by $T_{a}^{n}(y)>1$, there exist unique $z_{1},z_{2}\in[g_{n}(y),y]$ such that $T_{a}^{n}(z_{1})=\frac{2a-1}{1-a}$ and $T_{a}^{n}(z_{2})=1$. Similarly to the previous cases, $g_{n}(y),z_{1}]$ and $z_{2},y]$ are the maximal subintervals of $g_{n}(y),y]$ contained in $\text{dom}(T^{n+1}_{a})$, and there are no other ones. It holds $z_{1},y\in D_{n+1}$ and $g_{n}(y),z_{2}\in D_{n+1}$. We also have $g_{n+1}(z_{1})=g_{n}(y)$, $g_{n+1}(y)=z_{2}$.
@@
Furthermore, in either of the three cases, the function $T_{a}^{n+1}$ is continuous and strictly increasing on $g_{n+1}(y),y]$ as a composition of $T_{a}$ and $T_{a}^{n}$, which have these properties on $1,\frac{1}{1-a}]$ and $g_{n}(y),y]$, respectively. Then $\delta_{n}$ is constant on $g_{n+1}(y),y]$. In the third case, $T_{a}^{n+1}$ and $\delta_{n}$ also have these respective properties on $g_{n+1}(z_{1}),z_{1}]$ because on this interval $T_{a}^{n}$ does not exceed $\frac{2a-1}{1-a}$ and $T_{a}$ is continuous and strictly increasing on $0,\frac{2a-1}{1-a}]$.
b) The proof for $a=\frac{2}{3}$ is analogous. It suffices to replace throughout $g_{n}(y),y]$ by $g_{n}(y-),y)$ for all $y\in D_{n}\setminus\{3\}$ and use that $T_{2/3}^{n}(y-)=3$ instead of $T^{n}_{a}(y)=\frac{1}{1-a}$. ∎
@@
because the left-hand side is integer. This proves inequality (32) when $\sigma_{a}(x)=\infty$.
@@
We first assume that $g_{n}(y)\neq y$. Since the functions $z\mapsto L_{k}(z)$ for $0\leq k\leq n$ are constant on $g_{n}(y),y]$ by Lemma 12.a, inequality (32) follows instantly from the result in the previous case if we prove that there is an $x^{\prime}\in[g_{n}(y),y]$ such that $\sigma_{a}(x^{\prime})=\infty$. We will use that there exist periodic orbits of $T_{a}$ starting arbitrarily close to the left of $y$. Consider two cases. If $T_{a}^{m}(y)=\frac{2a-1}{1-a}$ for a unique $0\leq m\leq n-1$, then $T_{a}^{m}<\frac{2a-1}{1-a}$ on $g_{n}(y),y)$ by Lemma 12.a. By piecewise continuity of $T_{a}$, the equation $z=T_{a}^{t_{1}(z)}(z)$ admits infinitely many solutions, which accumulate to $\frac{2a-1}{1-a}$ from the left. Pick any solution $z^{\prime}\in[T_{a}^{m}(g_{n}(y)),\frac{2a-1}{1-a})$, and then take the unique $x^{\prime}\in[g_{n}(y),y]$ that solves $T_{a}^{m}(x^{\prime})=z^{\prime}$. In the second case when $y=\frac{1}{1-a}$, we argue similarly and take any $x^{\prime}\geq g_{n}(y)$ that satisfies $x^{\prime}=T_{a}^{t_{1}(x^{\prime})+1}(x^{\prime})$.
We now assume that $g_{n}(y)=y$. This is possible only when $G_{n}$ intersects $D_{n}$, that is when the orbit of zero hits $\frac{2a-1}{1-a}$. Then $T_{a}^{k_{1}}(x)=0$ and $T_{a}^{k_{2}}(x)=\frac{2a-1}{1-a}$ for the unique $k_{1}$, $k_{2}$ such that $0\leq k_{1}<k_{2}\leq n-1$. Thus, $x\in G_{k_{2}}$ and by Lemma 12.a, there exists a unique $y^{\prime}\in D_{k_{2}}$ such that $g_{k_{2}}(y^{\prime})=x$. In other words, $\varkappa_{a}(z)=k_{2}$ for every $z\in(g_{k_{2}}(y^{\prime}),y^{\prime})$. Clearly, $g_{k_{2}}(y^{\prime})\neq y^{\prime}$, and we can apply the result of the previous case for all $k$ satisfying $0\leq k\leq k_{2}$. Therefore, since the functions $z\mapsto L_{k}(z)$ for $0\leq k\leq k_{2}$ are constant on $g_{k_{2}}(y^{\prime}),y^{\prime}]$,